15. If a 10 kg block sits on a horizontal surface with a coefficient of uk=0.2. The block is at rest when a force of (4t) lb to the right is applied, where t is in seconds. What would be the velocity of the block after 2 seconds?
Answer #1

Net force on the block in horizontal direction = F-ukmg = 4*4.448t-0.2*9.8*10 = 17.793t-19.6 Newtons

Newton's second law. F = ma
a = 17.793t-19.6/10 = 1.7793t-1.96 m/s2
dv/dt = 1.7793t-1.96
v = 1.7793t^2/2 - 1.96t
since initial velocity = 0
at t = 2 s
v = 1.7793*4-1.96*2 = 3.1972 m/s